OpenStudy (anonymous):

Find the sum of the first 15 terms of the arithmetic sequence 4, 7, 10, 13, …

OpenStudy (anonymous):

+3

OpenStudy (anonymous):

so u got it?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

mmm r u talking about the diference it is 3

OpenStudy (firejay5):

I can explain

OpenStudy (firejay5):

How do you get 4 + ? = 7

OpenStudy (anonymous):

It was in the question

OpenStudy (anonymous):

well so i let it to u firejay

OpenStudy (firejay5):

@julian25 See if you can answer my question

OpenStudy (firejay5):

@DaWizjr Do you have a problem? Any questions

OpenStudy (anonymous):

first u need to find the diference do u know what it is dawzjr?

OpenStudy (firejay5):

4 + (3) = 7, 7 + (3) = 10, 10 + (3) = 13. The answer is in the parentheses.

OpenStudy (firejay5):

Make sense

OpenStudy (anonymous):

Well duh but its asking for the sum of the first 15 terms

OpenStudy (anonymous):

"sum"

OpenStudy (firejay5):

If you know that's is adding 3 each time, you can go on from 13 + 3 = 16, 16 is the forth term, you do the rest

OpenStudy (anonymous):

^ you are write but its asking for the "sum" so I brb

OpenStudy (firejay5):

Figure out what the 15 terms are and adding them together

OpenStudy (firejay5):

add all the terms together is what I meant

OpenStudy (anonymous):

u must know the diference in order to get the sum

OpenStudy (firejay5):

It's not the difference, it's the sum of 3

OpenStudy (firejay5):

Meaning it's adding 3 each time

OpenStudy (anonymous):

t1 = 2 t10 = 6.5 S10 = 42.5 t1 = 2 t20 = 11.5 S20 = 135

OpenStudy (anonymous):

THis is an example

OpenStudy (anonymous):

it is called common diference

OpenStudy (anonymous):

It says what is the sum of the first 10 terms of the sequence that starts with 2,2.5,3,3.5... what is the sum of the first 20 terms?

OpenStudy (anonymous):

t1 = 2 t10 = 6.5 S10 = 42.5 t1 = 2 t20 = 11.5 S20 = 135

OpenStudy (firejay5):

don't make it complicated

OpenStudy (anonymous):

Find the sum of the first 15 terms of the arithmetic sequence 4, 7, 10, 13, …

OpenStudy (anonymous):

4,7,10,13,16,19,21,24,27,30,33,36,39,41,44,47,49

OpenStudy (firejay5):

Incorrect, but right

OpenStudy (anonymous):

so 15=44

OpenStudy (anonymous):

!44!

OpenStudy (firejay5):

nope it's 46

OpenStudy (anonymous):

How?

OpenStudy (firejay5):

You added wrong 19 + 3 = 22 + 3 = 25, and so on

OpenStudy (anonymous):

ok cool